# The FinBen: An Holistic Financial Benchmark for Large Language Models

### v1

@@ Abstract
LLMs have transformed NLP and shown promise in various fields, yet their potential in finance is underexplored due to a lack of thorough evaluations and the complexity of financial tasks. This along with the rapid development of LLMs, highlights the urgent need for a systematic financial evaluation benchmark for LLMs. In this paper, we introduce FinBen, the first comprehensive open-sourced evaluation benchmark, specifically designed to thoroughly assess the capabilities of LLMs in the financial domain. FinBen encompasses 35 datasets across 23 financial tasks, organized into three spectrums of difficulty inspired by the Cattell-Horn-Carroll theory, to evaluate LLMs’ cognitive abilities in inductive reasoning, associative memory, quantitative reasoning, crystallized intelligence, and more. Our evaluation of 15 representative LLMs, including GPT-4, ChatGPT, and the latest Gemini, reveals insights into their strengths and limitations within the financial domain. The findings indicate that GPT-4 leads in quantification, extraction, numerical reasoning, and stock trading, while Gemini shines in generation and forecasting; however, both struggle with complex extraction and forecasting, showing a clear need for targeted enhancements. Instruction tuning boosts simple task performance but falls short in improving complex reasoning and forecasting abilities. FinBen seeks to continuously evaluate LLMs in finance, fostering AI development with regular updates of tasks and models [^1] https://github.com/The-FinAI/PIXIU.
The FinBen: An Holistic Financial Benchmark for Large Language Models
[…]
@@ 1 Introduction
Recently, Large Language Models (LLMs) Brown et al. (2020) such as ChatGPT [^2] https://openai.com/chatgpt and GPT-4 OpenAI (2023a), have reshaped the field of natural language processing (NLP) and exhibited remarkable capabilities in specialized domains across mathematics, coding, medicine, law, and finance Bubeck et al. (2023). With their increasing model size and extensive pre-training data, LLMs have developed the emergent capacity for in-context learning, enabling them to perform remarkably across a wide range of domain-specific tasks in zero-shot and few-shot settings Wei et al. (2023). Within the financial domain, recent several studies Xie et al. (2023a); Lopez-Lira and Tang (2023); Li et al. (2023b); Xie et al. (2023b) have shown the great potential of advanced LLMs such as GPT-4 on financial text analysis and prediction tasks. While their potential is evident, a comprehensive understanding of their capabilities and limitations for finance, remains largely unexplored. This is due to a lack of extensive evaluation studies and benchmarks, and the inherent complexities associated with the professional nature of financial tasks.
Existing financial domain evaluation benchmarks including FLUE Shah et al. (2022), BBT-CFLEB Lu et al. (2023), and PIXIU Xie et al. (2023b), have a limited scope and are solely focused on financial NLP tasks, primarily targeting language understanding abilities where LLMs have already been extensively evaluated. As shown in Table 1, they fail to capture other crucial facets of the financial domain, such as comprehending and extracting domain-specific financial knowledge and resolving realistic financial tasks. As such, their efficacy in evaluating and understanding LLM performance is limited.
Benchmark Datasets Quantification Extraction Understanding Generation Forecasting Trading
BloombegGPT-FLUE (Wu et al., 2023) 5 3 0 2 0 0 0
PIXIU (Xie et al., 2023b) 15 4 2 2 2 5 0
FINANCEBENCH (Islam et al., 2023) 1 0 0 1 0 0 0
BizBench (Koncel-Kedziorski et al., 2023) 9 2 1 5 1 0 0
CALM-bench (Feng et al., 2023) 9 0 0 0 0 9 0
BBT-CFLEB Lu et al. (2023) 6 2 2 1 1 0 0
CFBenchmark (Li et al., 2023a) 8 3 2 0 2 0 0
DISC-FinLLM (Chen et al., 2023c) 9 3 3 2 1 0 0
Fin-Eva (Team, 2023a) 1 1 0 0 0 0 0
FinBen 35 10 5 5 2 12 1
@@
Table 1: Comparison of different financial benchmarks based on the number of datasets used and their distribution across various tasks including quantification, extraction, understanding, generation, forecasting, and trading.
Furthermore, while there are newly released benchmarks in the general domain, such as MMLU Hendrycks et al. (2020), HELM Liang et al. (2022) and BIG-bench Srivastava et al. (2023) compiling massive tasks across numerous institutions, they do not extend to the financial domain. The fast progression of LLMs, coupled with an incomplete understanding of their abilities and behavior, highlights the need for a systematic financial evaluation benchmark dedicated to these models.
How should an effective systematic financial evaluation benchmark be designed? We believe it should fulfill the following criteria: 1) Broad coverage: It should cover a broad spectrum of tasks to capture the financial domain’s complexity, incorporating both linguistic understanding and diverse skills like knowledge extraction, text generation, and numerical reasoning et al. 2) Real-world application orientation: The benchmark should focus on real-world scenarios, including stock market analysis and trading, highlighting LLMs’ practical application capabilities. 3) Inclusion of financial domain-specific characteristics: It also needs to address the unique aspects of finance, embedding tasks that demand specific knowledge, terminology, and concepts, demonstrating LLMs’ proficiency in the field. 4) Consideration of human-level cognition: It should gauge human-like cognitive abilities, evaluating LLMs on decision-making, problem-solving, and abstract reasoning within financial contexts.
To bridge this gap, we propose FinBen, the first open-sourced [^3] We will release all resources to the research community. comprehensive evaluation benchmark designed for assessing the capabilities of LLMs in the financial domain. As shown in Figure 1, FinBen includes 35 datasets spanning 23 financial tasks organized into three Spectrums of difficulty inspired by the Cattell-Horn-Carroll (CHC) theory (Schneider and McGrew, 2012) in the fields of psychology and education, to assess LLMs across various cognitive domains, including inductive reasoning, associative memory, quantitative reasoning, crystallized intelligence, fluid intelligence, and general intelligence. Spectrum I is comprised of foundational tasks inclduing Quantification, Extraction, and Numerical Understanding, laying the groundwork for basic cognitive skills. Moving up, Spectrum II delves into more complex Generation and Forecasting tasks, demanding enhanced cognitive involvement. At the apex, Spectrum III focuses on the sophisticated stock trading task, exemplifying the application of General Intelligence.
Figure 1: Evaluation framework of FinBen.
In align with the above criteria, FinBen distinguishes from existing benchmarks from the breadth and depth of its coverage, as well as its uniquely tailored focus on the financial domain: 1) Wide coverage: FinBen integrates classic NLP tasks (text analysis, knowledge extraction, question answering) with finance-specific challenges (numeric labeling) and innovates by assessing LLMs on real-world financial applications (stock prediction, credit scoring) and for the first time directly assess the trading performance of LLMs. This broad approach unveils LLMs’ strengths and limitations in finance comprehensively. 2) Multi-data modality and diversity of text types: FinBen distinguishes itself by embracing diverse data forms and text types for its tasks, including news, tweets, earnings calls, financial documents, tables, and time-series data. This variety facilitates a thorough assessment of LLMs’ comprehension and generation of financial content, highlighting their real-world utility. 3) Diverse difficulty levels: the FinBen incorporates tasks of varying difficulty levels, from simpler fundamental tasks like news headline classification, to advenced congnitive engagement tasks such as the stock movement prediction, and even more complex general intelligence tasks that even challenge humans, such as stock trading. This range enables a nuanced evaluation of LLMs, fully mapping their strengths and weaknesses in finance.
We test 15 representative general LLMs such as GPT-4, ChatGPT and the latest Gemini, and financial LLMs in FinBen, and have following findings: 1) GPT-4 outperforms all others in quantification, extraction, numerical reasoning, and the intricate stock trading task, whereas Gemini excels in generation and forecasting tasks. 2) while state-of-the-art (SOTA) LLMs such as GPT-4 demonstrate superior capabilities in quantification and simple extraction tasks, they fall short in areas requiring advanced numerical reasoning and complex information extraction. Notably, these LLMs show promise in the demanding stock trading task, yet there is a pronounced need for improvement in text generation and forecasting tasks, which rely heavily on crystallized and fluid intelligence. 3) Instruction tuning is an effective way for improve the performance on quantification and simple extraction tasks, while it is less useful on other tasks such as numerical reasoning, generation and forecasting.
2 The FinBen
Our benchmark framework evaluates financial LLMs through a hierarchy inspired by the Cattell-Horn-Carroll (CHC) theory (Schneider and McGrew, 2012), defining cognitive abilities in three Spectrums. Spectrum I includes foundational tasks like Quantification (Inductive Reasoning) using classification tasks, Extraction (Associative Memory) covering information extraction tasks, and Numerical Understanding (Quantitative Reasoning) covering numerical reasoning tasks. Spectrum II advances to Generation (Crystallized Intelligence) covering generaltion task, and Forecasting (Fluid Intelligence) with prediction tasks, requiring deeper cognitive engagement. The pinnacle, Spectrum III, encompasses strategic decision-making in trading using the current state-of-art (SOTA) financial LLM agent Yu et al. (2023) with the stock trading task, showcasing General Intelligence (McGrew, 2009). This structured approach allows for nuanced assessment of LLMs’ financial analytical capabilities across varied cognitive demands. Table 2 and Figure 2 shows all tasks, datasets, data statistics and evaluation metrics covered by FinBen [^4] For detail instructions of each dataset, please see Appendix C.
Figure 2: Evaluation datasets of FinBen.
Data Task Valid Test Evaluation License
FPB (Malo et al., 2014) sentiment analysis 775 970 F1, Accuracy CC BY-SA 3.0
FiQA-SA (Maia et al., 2018a) sentiment analysis 188 235 F1 Public
TSA Cortis et al. (2017) sentiment analysis - 561 F1, Accuracy CC BY-NC-SA 4.0
Headlines (Sinha and Khandait, 2021) news headline classification 1,141 2,283 Avg F1 CC BY-SA 3.0
FOMC Shah et al. (2023a) hawkish-dovish classification - 496 F1, Accuracy CC BY-NC 4.0
FinArg-ACC Sy et al. (2023) argument unit classification - 969 F1, Accuracy CC BY-NC-SA 4.0
FinArg-ARC Sy et al. (2023) argument relation classification - 496 F1, Accuracy CC BY-NC-SA 4.0
MultiFin Jørgensen et al. (2023) multi-class classification - 690 F1, Accuracy Public
MA Yang et al. (2020a) deal completeness classification - 500 accuracy,F1 Public
MLESG Chen et al. (2023a) ESG Issue Identification - 300 accuracy,F1 CC BY-NC-ND
NER (Alvarado et al., 2015) named entity recognition 103 980 Entity F1 CC BY-SA 3.0
FiNER-ORD (Shah et al., 2023b) named entity recognition - 1080 Entity F1 CC BY-NC 4.0
FinRED Sharma et al. (2022) relation extraction - 1,068 F1, Entity F1 Public
SC Mariko et al. (2020) causal classification - 8,630 F1,Entity F1 CC BY 4.0
CD Mariko et al. (2020) causal detection - 226 F1,Entity F1 CC BY 4.0
FinQA (Chen et al., 2021) question answering 883 1,147 EM Accuracy MIT License
TATQA Zhu et al. (2021) question answering - 1,668 F1,EM Accuracy MIT License
ConvFinQA (Chen et al., 2022) question answering 2,210 1,490 EM Accuracy MIT License
FNXL Sharma et al. (2023) numeric labeling - 318 F1,EM Accuracy Public
FSRL Lamm et al. (2018) token classification - 97 F1, EM Accuracy MIT License
ECTSum Mukherjee et al. (2022) text summarization - 495 ROUGE, BERTScore, BARTScore Public
EDTSum Zhou et al. (2021) text summarization - 2000 ROUGE, BERTScore, BARTScore Public
BigData22 (Soun et al., 2022) stock movement prediction 798 1,470 Accuracy, MCC Public
ACL18 (Xu and Cohen, 2018) stock movement prediction 2,560 3,720 Accuracy, MCC MIT License
CIKM18 (Wu et al., 2018) stock movement prediction 431 1,140 Accuracy, MCC Public
German Hofmann (1994) credit scoring - 1000 F1, MCC CC BY 4.0
Australian Quinlan credit scoring - 690 F1, MCC CC BY 4.0
LendingClub Feng et al. (2023) credit scoring 1,344 2,690 F1, MCC CC0 1.0
ccf Feng et al. (2023) fraud detection 1,138 2,278 F1, MCC (DbCL) v1.0
ccfraud Feng et al. (2023) fraud detection 1,047 2,097 F1, MCC Public
polish Feng et al. (2023) financial distress identification 868 1,736 F1, MCC CC BY 4.0
taiwan Feng et al. (2023) financial distress odentification 681 1,364 F1, MCC CC BY 4.0
ProtoSeguro Feng et al. (2023) claim analysis 1,189 2,381 F1, MCC Public
travelinsurance Feng et al. (2023) claim analysis - 3,800 F1, MCC (ODbL) v1.0
fintrade Yu et al. (2023) stock trading - 3,384 CR, SR, DV, AV, MD MIT License
@@ 2 The FinBen
Table 2: The tasks, datasets, data statistics, and evaluation metrics included in FinBen.
2.1 Spectrum I: Foundamental Tasks
Spectrum I including 20 datasets from 16 tasks to evalaute financial LLMs from three perspectives including Quantification (Inductive Reasoning), Extraction (Associative Memory) and Numerical Understanding (Quantitative Reasoning).
Quantification. The auantification task include 8 classification tasks for evaluating financial LLMs, to transform financial text into categorical labels and numerical scores. As inductive reasoning (Ir), it requires LLMs to discern patterns and quantify sentiments within financial narratives. 1) Sentiment analysis focuses on extracting sentiment information from financial texts. We utilize two datasets: the Financial Phrase Bank (FPB) (Malo et al., 2014), FiQA-SA (Maia et al., 2018a), and TSA Cortis et al. (2017) dataset. 2) News headline classification analyzes additional information, like price movements in financial texts, using the Headlines dataset (Sinha and Khandait, 2021), which includes news about "gold" from 2000 to 2019 and their 9 corresponding tags. 3) Hawkish-Dovish classification aims to classify sentences from monetary policy texts as ’hawkish’ or ’dovish,’ focusing on the nuanced language and economic implications of financial texts, using the FOMC Shah et al. (2023a) dataset. 4) Argument unit classification categorizes sentences as claims or premises using the FinArg AUC dataset Sy et al. (2023). 5) Argument relation detection identifies relationships (attack, support, or irrelevant) between social media posts using the FinArg ARC dataset Sy et al. (2023). 6) Multi-class classification targets categorizing a variety of financial texts, including analyst reports, news articles, and investor comments, utilizing the MultiFin dataset  Jørgensen et al. (2023). 7) Deal completeness classification predicts if mergers and acquisitions events are "completed" or remain "rumors" based on news and tweets, employing the MA dataset Yang et al. (2020a). 8) ESG issue identification focuses on detecting Environmental, Social, and Governance (ESG) concerns in financial documents using the MLESG dataset Chen et al. (2023a). For all datasets, evaluation utilizes the accuracy and F1 Score.
@@ 2.1 Spectrum I: Foundamental Tasks
Extraction. The extraction task including 5 datasets from 4 information extraction tasks, evaluating LLMs’ ability to accurately retrieve specific financial information from large datasets, a process tied closely to Associative Memory (Ma). 1) Named entity recognition extracts entities like LOCATION, ORGANIZATION, and PERSON from financial agreements and SEC filings, using the NER Alvarado et al. (2015) and FINER-ORD Shah et al. (2023b) datasets. 2) Relation extraction identifies relationships such as "product/material produced" and "manufacturer" in financial news and earnings transcripts with the FINRED dataset Sharma et al. (2022). 3) Causal classification discerns whether sentences from financial news and SEC filings convey causality using the SC dataset Mariko et al. (2020). 4) Causal detection identifies cause and effect spans in financial texts with the CD dataset Mariko et al. (2020). The evaluation of these tasks is focused on the F1 score Goutte and Gaussier (2005) and Entity F1 score Derczynski (2016).
Understanding. The understanding task includes 5 datasets from 4 numerical reasoning tasks, challenging LLMs to interpret and analyze complex numerical data and intricate financial statistics, associted with the Quantitative Reasoning (Gq) ability. 1) Question answering focuses on solving questions through multi-step numerical reasoning with financial reports and tables, utilizing the FinQA Chen et al. (2021) and TATQA Zhu et al. (2021) dataset. 2) Multi-turn question answering is an extension on QA with multi-turn questions and answers based on financial earnings reports and tables, using the ConvFinQA dataset Chen et al. (2022). 3) Numeric labeling aims at tagging numeric spans in financial documents using 2,794 labels with the FNXL dataset Sharma et al. (2023). 4) Token classification aims at identifying common attributes and comparative elements in textual analogies by extracting analogy frames, utilizing the FSRL dataset Lamm et al. (2018). Entity F1 score Derczynski (2016) and the Exact Match Accuracy (EMAcc) metric Kim et al. (2023) are used to evaluate these tasks.
2.2 Spectrum II: Advanced Cognitive Engagement
Spectrum II has 14 datasets across 6 tasks designed to assess the Generation (Crystallized Intelligence) and Forecasting (Fluid Intelligence) capabilities of LLMs, requiring deeper cognitive engagement.

### v2

# FinBen: An Holistic Financial Benchmark for Large Language Models

@@ Abstract
LLMs have transformed NLP and shown promise in various fields, yet their potential in finance is underexplored due to a lack of comprehensive evaluation benchmarks, the rapid development of LLMs, and the complexity of financial tasks. In this paper, we introduce FinBen, the first extensive open-source evaluation benchmark, including 36 datasets spanning 24 financial tasks, covering seven critical aspects: information extraction (IE), textual analysis, question answering (QA), text generation, risk management, forecasting, and decision-making. FinBen offers several key innovations: a broader range of tasks and datasets, the first evaluation of stock trading, novel agent and Retrieval-Augmented Generation (RAG) evaluation, and three novel open-source evaluation datasets for text summarization, question answering, and stock trading. Our evaluation of 15 representative LLMs, including GPT-4, ChatGPT, and the latest Gemini, reveals several key findings: While LLMs excel in IE and textual analysis, they struggle with advanced reasoning and complex tasks like text generation and forecasting. GPT-4 excels in IE and stock trading, while Gemini is better at text generation and forecasting. Instruction-tuned LLMs improve textual analysis but offer limited benefits for complex tasks such as QA. FinBen has been used to host the first financial LLMs shared task at the FinNLP-AgentScen workshop during IJCAI-2024, attracting 12 teams. Their novel solutions outperformed GPT-4, showcasing FinBen’s potential to drive innovation in financial LLMs. All datasets, results, and codes are released for the research community [^1] https://github.com/The-FinAI/PIXIU.
@@ 1 Introduction
Recently, Large Language Models (LLMs) (Brown et al., 2020) such as ChatGPT [^2] https://openai.com/chatgpt and GPT-4 (OpenAI, 2023a), have reshaped the field of natural language processing (NLP) and exhibited remarkable capabilities in specialized domains across mathematics, coding, medicine, law, and finance (Bubeck et al., 2023). Within the financial domain, recent several studies (Xie et al., 2023a; Lopez-Lira and Tang, 2023; Li et al., 2023c; Xie et al., 2023b) have shown the great potential of advanced LLMs such as GPT-4 on financial text analysis and prediction tasks. While their potential is evident, a comprehensive understanding of their capabilities and limitations for finance remains largely unexplored. This is due to a lack of extensive evaluation studies and benchmarks, and the inherent complexities associated with the professional nature of financial tasks.
Existing financial domain evaluation benchmarks including PIXIU (Xie et al., 2023b), FinanceBench (Islam et al., 2023) and BizBench (Koncel-Kedziorski et al., 2023), have limited Evaluation Tasks and primarily focus on Financial NLP Tasks (As shown in Table 1). Most existing benchmarks cover only a small number of evaluation tasks and are centered on evaluating NLP capabilities, such as information extraction and QA. While PIXIU stands out by covering the highest number of tasks, it includes only one evaluation task in most categories. This narrow focus limits their ability to comprehensively evaluate LLMs across the diverse and complex landscape of financial applications, such as forecasting, risk management, and decision-making. It is insufficient for a thorough evaluation of LLM capabilities, especially in the financial area.
@@
Table 1: Comparison of different financial benchmarks based on the number of tasks and datasets used, and the task number distribution across various aspects including information extraction (IE), textual analysis (TA), question answering (QA), text generation (TG), risk management (RM), forecasting (FO), and decision-making (DM).
Benchmark Language Dataset Task IE TA QA TG RM FO DM
CFBenchmark Chinese 8 7 1 3 3
Fin-Eva Chinese 1 1 1
PIXIU English 15 8 1 3 1 1 1 1
FinanceBench English 1 1 1
BizBench English 8 5 2 2 1
FinBen English 36 24 6 8 3 1 4 1 1
Figure 1: Evaluation datasets for FinBen, with colors indicating dataset sizes ranging from 100 to 4,000.
To bridge this gap, we propose FinBen, a novel comprehensive open-source evaluation benchmark developed through the collaborative efforts of experts in both computer science and finance. FinBen comprises 36 datasets spanning 24 financial tasks, meticulously organized to assess LLMs across seven critical aspects: information extraction (IE), textual analysis (TA), question answering (QA), text generation (TG), risk management (RM), forecasting (FO), and decision-making (DM). Each category targets specific skills of financial data handling and analysis, ensuring a thorough evaluation of LLMs and showcasing their proficiency in managing complex financial scenarios.
FinBen introduces several innovations compared to existing benchmarks: 1) New Tasks: FinBen introduces a significantly larger number of tasks and datasets, making it the most holistic benchmark for financial LLMs with the highest number of tasks and datasets. This extensive range provides a more robust evaluation of LLM capabilities in diverse financial contexts. 2) Broader Coverage: Covering seven aspects of the financial sector, FinBen is the first benchmark to include the evaluation of stock trading, which is the fundamental task in the financial sector, involving complex decision-making processes that impact market dynamics and investment strategies. 3) New Evaluation Strategy: FinBen is the first benchmark to include agent-based evaluation and Retrieval-Augmented Generation (RAG) based evaluation. These innovative strategies provide a more dynamic and realistic assessment of LLMs, reflecting their ability to interact with and retrieve relevant information from vast datasets. 4) Novel Datasets: FinBen proposes three novel open-source datasets of text summarization, QA, and stock trading tasks for the research community, pushing the boundaries of what LLMs can achieve and setting a new standard for dataset comprehensiveness. 5) Empowering Financial LLMs Research: Based on FinBen, we hosted the first shared task (see Appendix F for more details) focused on financial LLMs at the FinNLP-AgentScen workshop during IJCAI-2024  [^3] https://sites.google.com/nlg.csie.ntu.edu.tw/finnlp-agentscen. This event attracted 12 teams, leveraging our benchmark to develop novel LLMs-based solutions within the financial domain. Remarkably, the proposed methods achieved superior performance compared to GPT-4, demonstrating the benchmark’s potential to foster innovation and advance the state-of-the-art (SOTA) in financial LLMs.
Based on FinBen, we access 15 representative general LLMs such as GPT-4, ChatGPT, and the latest Gemini, and financial LLMs, and have the following findings: 1) Superior Capabilities with Limitations: While LLMs exhibit exceptional prowess in IE and textual analysis tasks, they underperform in areas necessitating advanced reasoning and complex IE, such as text generation and forecasting. 2) Potential in Stock Trading: SOTA LLMs have demonstrated considerable promise in stock trading applications. However, there remains significant room for improvement due to their limitations in reasoning and comprehensive forecasting abilities. 3) Closed-Source Superiority: Closed-source commercial LLMs continue to lead in performance within the financial domain. Specifically, GPT-4 excels in IE, text analysis, QA, and intricate stock trading tasks, while Gemini shows superior capabilities in text generation and forecasting. 4) Open-Source Improvements and Limitations: While open-source, instruction-tuned financial LLMs have shown notable enhancements in textual analysis and IE tasks, the advantages of instruction-tuning are less pronounced when it comes to complex tasks such as QA, text generation, and forecasting.
In summary, the main contributions of this paper are: 1) we propose FinBen, the first comprehensive open-sourced evaluation benchmark for LLMs in the financial domain, 2) we propose a novel taxonomy covering seven aspects for organizing financial evaluation tasks, 3) we propose three novel evaluation datasets for the research community, and 4) we conduct systematic evaluation of 15 LLMs using FinBen, showcasing their limitations and advantages, and highlighting directions for future work.
2 FinBen
In this section, we delve into the specifics of FinBen, detailing the evaluation taxonomy, data sources, and evaluation tasks.
2.1 The Taxonomy of Financial Evaluation Tasks
In the dynamic landscape of financial technology, evaluating the capabilities of LLMs necessitates a comprehensive and structured approach. We propose a novel taxonomy for financial evaluation tasks, categorizing and assessing LLMs across seven financial domains inspired by established taxonomies in financial tasks Cao (2022); Li et al. (2023b); Zhao et al. (2024): Information Extraction (IE), Textual Analysis (TA), Question Answering (QA), Text Generation (TG), Risk Management (RM), Forecasting (FO), and Decision-Making (DM). Information Extraction focuses on identifying key entities and relationships within financial documents, transforming unstructured data into structured insights Costantino and Coletti (2008). Textual Analysis delves into content and sentiment analysis of financial texts, aiding in market trend understanding Loughran and McDonald (2020). Question Answering evaluates the model’s ability to comprehend and respond to financial queries Maia et al. (2018). Text Generation assesses the production of coherent financial text La Quatra and Cagliero (2020). Risk Management involves evaluating creditworthiness, detecting fraud, and ensuring regulatory compliance Aziz and Dowling (2019). Forecasting predicts future financial trends, enabling strategic responses to market dynamics Abu-Mostafa and Atiya (1996). Finally, Decision-Making assesses the model’s proficiency in making informed financial decisions, such as developing trading strategies and optimizing investment portfolios Paiva et al. (2019).
2.2 Data Sources
FinBen’s evaluation tasks are drawn from three primary data sources: 1) open-sourced datasets from existing studies comprising open-sourced datasets originally released for non-LLM evaluation settings. Domain experts have designed diverse prompts and reformulated these datasets into instruction-response pairs, making them suitable for evaluating the zero-shot performance of LLMs. 2) datasets from existing evaluation benchmarks including datasets from existing evaluation benchmarks such as PIXIU. These datasets have already been transformed into the instruction tuning format, allowing for seamless integration and direct use in FinBen. 3) novel datasets introduced in this paper. These datasets are designed to address gaps in existing benchmarks and provide unique challenges for financial LLMs evaluation. Novel datasets include (As shown in Table 2): EDTSum: We develop the EDTSum dataset for the text summarization task, focusing on financial news. This dataset contains 2,000 news with their headlines as the ground-truth summarization manually selected and cleaned from Zhou et al. (2021). This dataset represents a significant contribution to evaluating the ability of LLMs to generate concise and informative summaries. FinTrade: We develop the FinTrade dataset specifically for the stock trading task, incorporating historical stock prices, news, and sentiment data. This dataset is the collection of 10 stocks over one year. Our contribution enables the assessments of LLMs in agent-based financial trading scenarios. Regulations: We develop the regulations dataset for the long-form QA task related to Over-the-Counter (OTC) derivatives and general financial matters within the European Union. It precisely maps these inquiries to relevant articles from the European Market Infrastructure Regulation (EMIR) and other directives, with updates ranging from 2013 to 2021. It tests LLMs’ ability to interpret and apply complex regulatory frameworks in real-world scenarios.
2.3 Tasks
@@
Table 2 and Figure 1 shows all tasks, datasets, data statistics, and evaluation metrics covered by FinBen [^4] For detailed instructions of each dataset, please see Appendix C.
Information extraction:
@@
It spans seven datasets across six information extraction tasks. 1) Named entity recognition extracts entities like LOCATION, ORGANIZATION, and PERSON from financial agreements and SEC filings, using the NER (Alvarado et al., 2015) and FINER-ORD (Shah et al., 2023b) datasets. 2) Relation extraction identifies relationships such as "product/material produced" and "manufacturer" in financial news and earnings transcripts with the FINRED dataset (Sharma et al., 2022). 3) Causal classification discerns whether sentences from financial news and SEC filings convey causality using the SC dataset (Mariko et al., 2020). 4) Causal detection identifies cause and effect spans in financial texts with the CD dataset (Mariko et al., 2020). 5) Numeric labeling tags numeric spans in financial documents using the FNXL dataset Sharma et al. (2023), focusing on automating the assignment of labels from a large taxonomy to numeral spans in sentences. 6) Textual analogy parsing involves identifying common attributes and comparative elements in textual analogies by extracting analogy frames, utilizing the FSRL dataset Lamm et al. (2018), which maps analogous facts to semantic role representations and identifies the analogical relations between them. The evaluation of these tasks is focused on the F1 score (Goutte and Gaussier, 2005), Entity F1 score (Derczynski, 2016), and the Exact Match Accuracy (EMAcc) metric (Kim et al., 2023).
Table 2: The tasks, datasets, data statistics, and evaluation metrics included in FinBen. We use only test data for evaluation. Datasets marked with an asterisk (*) are newly constructed by us, comprising 10.32% of the total data. EM Accuracy means the exact match accuracy.
Data Task Test Evaluation License
NER (Alvarado et al., 2015) named entity recognition 980 Entity F1 CC BY-SA 3.0
FiNER-ORD (Shah et al., 2023b) named entity recognition 1080 Entity F1 CC BY-NC 4.0
FinRED (Sharma et al., 2022) relation extraction 1,068 F1, Entity F1 Public
SC (Mariko et al., 2020) causal classification 8,630 F1,Entity F1 CC BY 4.0
CD (Mariko et al., 2020) causal detection 226 F1,Entity F1 CC BY 4.0
FNXL (Sharma et al., 2023) numeric labeling 318 F1,EM Accuracy Public
FSRL (Lamm et al., 2018) textual analogy parsing 97 F1, EM Accuracy MIT License
FPB (Malo et al., 2014) sentiment analysis 970 F1, Accuracy CC BY-SA 3.0
FiQA-SA (Maia et al., 2018) sentiment analysis 235 F1 Public
TSA (Cortis et al., 2017) sentiment analysis 561 F1, Accuracy CC BY-NC-SA 4.0
Headlines (Sinha and Khandait, 2021) news headline classification 2,283 Avg F1 CC BY-SA 3.0
FOMC (Shah et al., 2023a) hawkish-dovish classification 496 F1, Accuracy CC BY-NC 4.0
FinArg-ACC (Sy et al., 2023) argument unit classification 969 F1, Accuracy CC BY-NC-SA 4.0
FinArg-ARC (Sy et al., 2023) argument relation classification 496 F1, Accuracy CC BY-NC-SA 4.0
MultiFin (Jørgensen et al., 2023) multi-class classification 690 F1, Accuracy Public
MA (Yang et al., 2020a) deal completeness classification 500 F1, Accuracy Public
MLESG (Chen et al., 2023a) ESG Issue Identification 300 F1, Accuracy CC BY-NC-ND
FinQA (Chen et al., 2021) question answering 1,147 EM Accuracy MIT License
TATQA (Zhu et al., 2021) question answering 1,668 F1, EM Accuracy MIT License
*Regulations long-form question answering 254 ROUGE, BERTScore Public
ConvFinQA (Chen et al., 2022b) multi-turn question answering 1,490 EM Accuracy MIT License
ECTSum (Mukherjee et al., 2022) text summarization 495 ROUGE, BERTScore, BARTScore Public
*EDTSum text summarization 2000 ROUGE, BERTScore, BARTScore Public
BigData22 (Soun et al., 2022) stock movement prediction 1,470 Accuracy, MCC Public
ACL18 (Xu and Cohen, 2018) stock movement prediction 3,720 Accuracy, MCC MIT License
CIKM18 (Wu et al., 2018) stock movement prediction 1,140 Accuracy, MCC Public
German (Hofmann, 1994) credit scoring 1000 F1, MCC CC BY 4.0
Australian (Quinlan, [n. d.]) credit scoring 690 F1, MCC CC BY 4.0
LendingClub (Feng et al., 2023) credit scoring 2,690 F1, MCC CC0 1.0
ccf (Feng et al., 2023) fraud detection 2,278 F1, MCC (DbCL) v1.0
ccfraud (Feng et al., 2023) fraud detection 2,097 F1, MCC Public
polish (Feng et al., 2023) financial distress identification 1,736 F1, MCC CC BY 4.0
taiwan (Feng et al., 2023) financial distress identification 1,364 F1, MCC CC BY 4.0
ProtoSeguro (Feng et al., 2023) claim analysis 2,381 F1, MCC Public
travelinsurance (Feng et al., 2023) claim analysis 3,800 F1, MCC (ODbL) v1.0
*FinTrade stock trading 3,384 CR, SR, DV, AV, MD MIT License
Textual analysis:
This encompasses eight classification tasks for evaluating LLMs. 1) Sentiment analysis focuses on extracting sentiment information (positive, negative, or neutral) from financial texts, using three datasets: the Financial Phrase Bank (FPB) (Malo et al., 2014), FiQA-SA (Maia et al., 2018), and TSA (Cortis et al., 2017) dataset. 2) News headline classification analyzes additional information, like price movements in financial texts, using the Headlines dataset (Sinha and Khandait, 2021). 3) Hawkish-Dovish classification aims to classify sentences from monetary policy texts as ’hawkish’ or ’dovish’ focusing on the nuanced language and economic implications of financial texts, using the FOMC (Shah et al., 2023a) dataset. 4) Argument unit classification categorizes sentences as claims or premises using the FinArg AUC dataset (Sy et al., 2023). 5) Argument relation detection identifies relationships (attack, support, or irrelevant) between social media posts using the FinArg ARC dataset (Sy et al., 2023). 6) Multi-class classification targets categorizing a variety of financial texts, including analyst reports, news articles, and investor comments, utilizing the MultiFin dataset (Jørgensen et al., 2023). 7) Deal completeness classification predicts if mergers and acquisitions events are "completed" or remain "rumors" based on news and tweets, employing the MA dataset (Yang et al., 2020a). 8) ESG issue identification focuses on detecting Environmental, Social, and Governance (ESG) concerns in financial documents using the MLESG dataset (Chen et al., 2023a). For all datasets, evaluation utilizes the accuracy and F1 Score.
Question answering.
It includes 3 datasets from 2 QA tasks, challenging LLMs to respond to financial queries. 1) QA focuses on solving questions through multi-step numerical reasoning with financial reports and tables, utilizing the FinQA (Chen et al., 2021) and TATQA (Zhu et al., 2021) dataset. 2) Multi-turn QA is an extension of QA with multi-turn questions and answers based on financial earnings reports and tables, using the ConvFinQA dataset (Chen et al., 2022b). F1 score (Derczynski, 2016) and the Exact Match Accuracy (EMAcc) metric (Kim et al., 2023) are used to evaluate these tasks. 2) Long-form QA involve presenting models with complex, detailed questions that require extensive and nuanced answers, often incorporating legal interpretations and practical applications. In our evaluation, we utilize our newly proposed Regulations dataset, which focuses on intricate questions and answers related to financial regulations like EMIR. We assess the model responses using ROUGE (Lin, 2004) and BERTScore (Zhang et al., 2019).